# Should we pre-train a decoder in contrastive learning for dense prediction tasks?

### v1

@@ 3.4 Evaluation protocol
We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks. We either transfer the pre-trained encoder or both the encoder and decoder for each fine-tuning experiment. All fine-tuning results reported in this paper are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware. We perform object detection and instance segmentation using Detectron2 package [47] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
Object detection and instance segmentation: We perform object detection on Pascal VOC dataset [18] and object detection and instance segmentation on COCO dataset [31]. We fine-tune a Faster R-CNN with a FPN C4 backbone using VOC trainval07+12 dataset for 24000 iterations and evaluate on the VOC test2007 set following previous works [22, 50]. For COCO object detection and instance segmentation, we fine-tune a Mask R-CNN with FPN backbone on the COCO train2017 dataset with 90000 iterations and evaluate on COCO val2017 split following SlotCon [46]. We initialize the network either with a pre-trained ResNet-50 encoder or with both a ResNet-50 encoder and a pre-trained FPN decoder, and fine-tune end-to-end in all cases. All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB memory.
Semantic segmentation on in-domain datasets: We perform semantic segmentation on Pascal VOC [18] and Cityscapes [13] datasets that contain scene-centric data following SlotCon [46] and PixCon [37]. A ResNet-50 encoder with a two-layer FCN decoder, similar to the FCN decoder used during pre-training, is used. For Pascal VOC, we train on VOC train_aug2012 set for 30000 iterations with a batch size of 16 and evaluate on VOC val2012 set. For Cityscapes, we fine tune on the train_fine set for 90000 iterations with a batch size of 16 and evaluate on the val_fine set. We initialize the network either with a pre-trained ResNet-50 encoder or with both the ResNet-50 encoder and the FCN decoder pre-trained and fine-tune end-to-end in all cases. All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB (Cityscapes) or 40GB (Pascal VOC) memory.
Semantic segmentation on out-of-domain datasets: We also evaluate generalizability to out-of-domain datasets for downstream tasks on different dataset sizes, considering 5, 25, and 100% randomly selected training samples. We fine-tune a ResNet-50 encoder coupled with an FCN decoder for semantic segmentation on REFUGE [36] and ISIC 2017 [10] dataset. REFUGE dataset contains 1200 retinal images in total representing 3 different classes: optic disk, optic cup, and background, divided into 400 training, 400 validation, and 400 testing samples. We train the ResNet-50-FCN network end-to-end for 80000 iterations with a batch size of 16 on the training set using one NVIDIA A100 GPU (40GB). The learning rate starts at 0.01 and decreases with a polynomial scheduler with a power of 0.9. We evaluate the model on the test set using the last iteration checkpoint and report the results.
ISIC 2017 is a skin-lesion segmentation dataset which consists of 2000 training, 150 validation, and 600 testing images. We fine-tune the same ResNet-50-FCN architecture on the training set for 24000 iterations on one NVIDIA V100 with 32GB of memory. Hyper parameters are the same as for REFUGE. We evaluate on the validation set every 500 iterations, select the best checkpoint and report the results of this checkpoint on the test set.
Table 1: Performance of DeCon-SL adaptation ($\alpha=0.5$) of SlotCon and DenseCL using FPN and FCN as options for decoders as well as baseline approaches. COCO dataset was used for pre-training. The same decoder architecture was used for pre-training and fine-tuning. The architecture was then fine-tuned on various segmentation and object detection tasks. All fine-tuning results were averaged over three runs. Up and down arrows indicate better and worst performance compared to baseline, respectively.
Framework Adaptation Decoder Loss Transfer Object Detection Object Detection Instance Seg. Semantic Seg.
VOC COCO COCO VOC City
$L_{enc}$ $L_{dec}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
SlotCon None None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
DeCon-SL FPN 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$ - -
55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$ - -
FCN - - - - - - - - - 72.42  $\uparrow$ 75.67  $\downarrow$
- - - - - - - - - 72.58  $\uparrow$ 76.00  $\uparrow$
DenseCL None None 54.50 80.47 59.97 39.60 59.30 43.29 35.65 56.49 38.25 71.11 75.87
DeCon-SL FPN 54.71  $\uparrow$ 80.65  $\uparrow$ 60.44  $\uparrow$ 39.60 59.36  $\uparrow$ 43.31  $\uparrow$ 35.74  $\uparrow$ 56.45  $\downarrow$ 38.40  $\uparrow$ - -
54.41  $\downarrow$ 80.24  $\downarrow$ 60.01  $\uparrow$ 39.55  $\downarrow$ 59.41  $\uparrow$ 43.15  $\downarrow$ 35.71  $\uparrow$ 56.44  $\downarrow$ 38.34  $\uparrow$ - -
FCN - - - - - - - - - 71.31  $\uparrow$ 76.04  $\uparrow$
- - - - - - - - - 71.59  $\uparrow$ 76.17  $\uparrow$
@@ 4 Results
DeCon-SL: Table 1 shows the results of DeCon-SL adaptation. DeCon pre-trained components outperformed baseline models in downstream object detection and segmentation when only the encoder was transferred, except for Cityscapes semantic segmentation with SlotCon. The transfer of both a pre-trained encoder and an FCN decoder always improved the performance further compared to transferring the pre-trained encoder only. Transferring the encoder and the FPN decoder pre-trained with DeCon-SL adaptation of DenseCL did not always result in improving model performance for downstream tasks.
DeCon-ML: Table 2 shows the results of DeCon-ML adaptation. We can see that adding the decoder deep supervision and a channel dropout probability of 0.5 during pre-training always enhanced fine-tuning performance compared to baseline approach (SlotCon) and DeCon-SL adaptation in all object detection and instance segmentation tasks.
Table 2: Performance of DeCon-ML adaptation of SlotCon with $\alpha=0.5$ and $dropout=0.5$ pre-trained on COCO dataset. We used an FPN as a DeCon decoder during pre-training and fine-tuning. We averaged all fine-tuning results over three runs. Arrows indicate performance gain or loss compared to baseline.
Frame work Adaptation Loss Transfer Object Detection Object Detection Instance Segmentation
VOC COCO COCO
$L_{enc}$ $L_{dec}$ $L_{dds}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
SlotCon None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
DeCon-SL 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$
55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$
DeCon-ML 55.60  $\uparrow$ 81.84  $\uparrow$ 61.83  $\uparrow$ 41.10  $\uparrow$ 61.37  $\uparrow$ 45.01  $\uparrow$ 37.05  $\uparrow$ 58.46  $\uparrow$ 39.87  $\uparrow$
55.57  $\uparrow$ 81.77  $\uparrow$ 61.76  $\uparrow$ 41.15  $\uparrow$ 61.38  $\uparrow$ 44.89  $\uparrow$ 37.05  $\uparrow$ 58.43  $\uparrow$ 39.88  $\uparrow$
Impact of the channel dropout: Impact of channel dropout is highlighted in Tab. 3. The decoder deep supervision on its own did not always improve downstream performance when only transferring the encoder in DeCon-ML. When transferring both encoder and decoder, downstream performance improved on Pascal VOC but not on COCO, which was the dataset used for pre-training. When the decoder deep supervision was coupled with channel dropout, performance improved compared to DeCon-SL. The channel dropout probability impacted the downstream performance differently when transferring only the encoder or both encoder and decoder. A channel dropout probability of 0.5 led to the best performance when transferring the encoder only. When transferring both encoder and decoder, a lower channel dropout probability of 0.25 led to the best performance on Pascal VOC object detection, and COCO instance segmentation. Comparing the performance of DeCon-ML adaptation with an encoder-only pre-training, we observed improvements in all scenarios regardless of dropout values.
Table 3: Impact of channel dropout probability on DeCon-ML. We pre-trained a DeCon-ML adaptation of SlotCon framework on COCO dataset with different channel dropout probabilities. We used an FPN as a DeCon decoder during pre-training and fine-tuning. We averaged all fine-tuning results over three runs.
Transfer SSL Loss $\alpha$ Dropout VOC Object Detection COCO Object Detection COCO Instance Seg.
$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
Enc SlotCon 1 0 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
DeCon-SL 0.5 0 55.57 81.75 61.76 40.94 60.97 44.82 36.92 58.03 39.80
0.5 55.20 81.56 61.33 40.90 61.23 44.64 36.91 58.30 39.67
DeCon-ML 0 55.52 81.67 62.01 40.85 60.95 44.67 36.83 58.12 39.55
0.25 55.54 81.88 61.10 40.95 61.13 44.88 36.97 58.22 39.76
0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87
0.75 55.29 81.64 61.36 41.02 61.27 44.96 37.05 58.33 39.85
Enc + Dec DeCon-SL 0.5 0 55.40 81.48 61.21 41.05 61.34 44.85 37.03 58.33 39.86
DeCon-ML 0 55.70 81.89 61.69 40.97 61.37 44.66 37.08 58.45 39.88
0.25 55.79 81.79 61.93 41.06 61.36 44.99 37.13 58.52 39.86
0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88
0.75 55.37 81.57 61.48 41.11 61.45 44.91 37.13 58.55 39.99
Impact of encoder loss weight in DeCon: As shown in Tab. 4, the encoder loss weight $\alpha$ played a crucial role in downstream performance. Decreasing $\alpha$ down to 0 for DeCon-ML pre-training and down to 0.25 for DeCon-SL pre-training increased the fine-tuning performance. This pattern was also observed when transferring both the encoder and the decoder. The downstream performance of DeCon adaptation was higher than that of encoder-only pre-training (baseline, $\alpha=1$) for object detection and instance segmentation tasks, whether transferring only the encoder or both the encoder and decoder. For Cityscapes semantic segmentation, no pattern was observed regarding the relationship between the value of $\alpha$ and model performance.
@@
Table 4: Impact of the encoder weight $\alpha$. We pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder ($dropout=0.5$) and a DeCon-SL adaptation of SlotCon with an FCN decoder on COCO dataset. We averaged all fine-tuning results over three runs.
Transfer $\alpha$ SSL strategy Object Detection Object Detection Instance Seg. SSL strategy Semantic Seg.
VOC COCO COCO VOC City
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
Enc 1 SlotCon 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 SlotCon 71.50 75.95
0.75 DeCon-ML-FPN 0.5 Dropout ($L_{enc}+L_{dds}$) 55.33 81.68 61.34 40.98 61.28 44.76 36.95 58.34 39.72 DeCon-SL-FCN ($L_{enc}+L_{ds})$ 71.98 76.18
0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87 72.42 75.62
0.25 55.69 81.98 62.05 41.05 61.19 44.85 37.08 58.29 39.70 73.01 76.21
0 55.70 81.91 62.07 41.18 61.38 44.91 37.12 58.35 39.94 72.17 75.71
Enc + Dec 0.75 55.46 81.51 61.35 40.96 61.10 44.87 36.90 58.07 39.73 72.80 76.06
0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88 72.58 75.97
0.25 55.67 81.91 61.62 41.17 61.41 44.87 37.13 58.46 39.76 72.96 76.28
0 55.61 81.83 61.60 41.21 61.39 44.99 37.11 58.41 39.92 71.61 75.38
Heterogeneous pre-training and fine-tuning decoders: All the previous fine-tuning results were obtained using the pre-trained encoder of a DeCon adaptation, where the pre-training decoder matched the fine-tuning decoder. However, Tab. 5 shows that even if we pre-trained a DeCon adaptation of SlotCon with an FCN decoder and transferred the encoder only for an FPN-based downstream task, we obtained better performance than a baseline SlotCon pre-training. It is also true for the opposite scenario, when FPN decoder was used for pre-training DeCon-ML, and FCN decoder was used for fine-tuning.
Table 5: Heterogeneous pre-training and fine-tuning decoders. We pre-trained a DeCon-SL adaptation of SlotCon with an FCN decoder and transferred the pre-trained encoder to FPN-based models for downstream tasks. Similarly, we pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder and transferred the pre-trained encoder to FCN-based models for downstream tasks. We averaged fine-tuning results over three runs.
Pretraining $\alpha$ Drop Loss Obj. Det. Obj. Det. Inst. Seg. Semantic Seg.
VOC COCO COCO VOC City
$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP AP mIoU mIoU
SlotCon 1 0 55.16 40.81 36.80 71.50 75.95
DeCon-ML-FPN 0 0.5 - - - 72.92 76.45
DeCon-SL-FCN 0.5 0 55.29 40.96 36.98 - -
Comparison with state-of-the-art: Table 6 shows that our top-performing DeCon-ML and DeCon-SL adaptations achieved better performance than state-of-the-art methods on COCO object detection and instance segmentation, and VOC semantic segmentation tasks. A more detailed comparison is available in Supplementary Material Section 2.
[…]
Method Obj. Det. Inst. Seg. Semantic Seg.
VOC COCO COCO VOC City
APb APb APm mIoU mIoU
SlotCon ‡ [46] 55.16 40.81 36.80 71.50 75.95
PixCon-SR † [37] 57.55 40.81 36.84 72.95 76.62
DeCon-ML-SlotCon-FPN [ours] 55.70 41.18 37.12 72.92 76.45
DeCon-SL-SlotCon-FCN [ours] - - - 73.01 76.21
@@
DeCon-SL performance in out-of-domain tasks: Table 7 depicts the performance of a model pre-trained on COCO to out-of-domain datasets. We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased. Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if we transferred only the encoder. We also observed that pre-trained models outperformed models trained with randomly initialized weights. This gap was more notable in REFUGE dataset (over 20%).
@@
Table 7: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights. We pre-trained SlotCon and its DeCon-SL adaptation with an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets. We fine-tuned on REFUGE and ISIC datasets under various dataset sizes. We averaged the fine-tuning mIoU results over three runs.

### v2

@@ 3.4 Evaluation protocol
We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks. Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments. We initialize the network with a pre-trained backbone and fine-tune end-to-end in all cases. All fine-tuning results reported in this paper are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware. We perform object detection and instance segmentation using Detectron2 package [49] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
Object detection and instance segmentation: We perform object detection and instance segmentation on COCO dataset [31]. We fine-tune a Mask R-CNN with FPN backbone following SlotCon [48].
Semantic segmentation: We evaluate the semantic segmentation performance of a pre-trained ResNet-50 encoder with an FCN decoder on Pascal VOC [18] and Cityscapes [13] following SlotCon [48] and PixCon [39].
Semantic segmentation on out-of-domain datasets: We assess generalization on REFUGE [38] and ISIC 2017 [10] datasets using the same ResNet-50-FCN architecture, fine-tuned on 5, 25, and 100% of the training data.
More details on the training setups and hyperparameters are provided in the Supplementary Material Section 2.
@@ 4 Results
4.1 Experiments
Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML Small and Large ($\alpha=0$ and $dropout=0.5$) adaptations of SlotCon trained with a ResNet-50 backbone. The architecture was pre-trained for 800 epochs on COCO and COCO+ and for 200 epochs on ImageNet-1K, and then fine-tuned on various segmentation and object detection tasks. Only the pre-trained encoder was transferred for the fine-tuning. Results for PixCon pre-trained on ImageNet-1K are not reported, as the corresponding checkpoint or evaluation was not made available in the original publication. All fine-tuning results were averaged over three runs. means that we reimplemented the framework, * means that the results are taken from the original paper.
Pret. Dataset Framework Pret. Dec. Object Detection Instance Segmentation Semantic Segmentation
COCO COCO VOC City
AP AP50 AP75 AP AP50 AP75 mIoU mIoU
COCO PixCon [39]* - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62
SlotCon [48] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21
DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21
DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45
COCO+ PixCon [39]* - 41.2 - - 37.1 - - 73.9 77.0
SlotCon [48] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43
DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65
DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00
ImageNet-1K SlotCon [48] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15
DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51
ResNet-50 backbone: Table 1 presents the results of fine-tuning a ResNet-50 backbone pre-trained with the two best pre-training frameworks for dense downstream tasks and DeCon. A more comprehensive comparison with additional previously published methods is provided in the Supplementary Material 4. DeCon-ML-L is consistently establishing a new SATO for all tasks, except when pre-training is performed on COCO dataset for Pascal VOC and Cityscapes semantic segmentation fine-tuning, where it reduces the gap between PixCon and SlotCon. This supports the fact that the DeCon adpatation is robust across datasets. Secondly, we can see that all of DeCon versions consistently improve upon the base SlotCon framework for any of the pre-training datasets used. Notably, DeCon-ML-S, whose parameter count is similar to SlotCon, exhibits better performance than SlotCon.
@@ 4.1 Experiments
Table 2: Performance of DeCon-ML-L ($\alpha=0$ and $dropout=0.5$) adaptation of SlotCon using a ConvNeXt-small encoder. The architecture was pre-trained for 800 epochs on COCO, and then fine-tuned on various segmentation and object detection tasks. All fine-tuning results were averaged over three runs.
Pret. Dataset Framework Pret. Dec. Obj. Det. Inst. Seg. Sem. Seg.
COCO COCO VOC
AP AP mIoU
COCO SlotCon - 44.07 39.67 73.24
DeCon-ML-L FPN 44.71 40.37 73.81
ConvNeXt backbone: Table 2 presents the results of the fine-tuning of models pre-trained with a ConvNeXt-small backbone using DeCon-ML-L and the original SlotCon. A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones. Notably, the performance of the model as well as the performance improvement appear larger than for a smaller ResNet-50 backbone.
Table 3: Performance of DeCon-SL ($\alpha=0.5$) adaptation of DenseCL and PixPro using a ResNet-50 encoder. The architectures were pre-trained for 800 epochs on COCO, and then fine-tuned on Pascal VOC and Cityscapes semantic segmentation. All fine-tuning results were averaged over three runs.
Pret. Dataset Framework Pret. Dec. Transfer Sem. Seg.
VOC City
mIoU mIoU
COCO DenseCL - Enc 71.11 75.87
DeCon-SL (DenseCL) FCN Enc 71.31 76.04
Enc + Dec 71.59 76.17
PixPro - Enc 72.13 75.84
DeCon-SL-(PixPro) FCN Enc 72.18 76.16
Enc + Dec 73.08 76.19
Generalizability across frameworks: Table 3 presents semantic segmentation fine-tuning results of models pre-trained with DenseCL and PixPro, and their DeCon-SL adaptation. The performance of the model pre-trained with a decoder is improved compared to the base framework and improves even further when the pre-trained decoder is transferred, supporting the hypothesis that DeCon could be adapted to multiple frameworks. Transferring the pre-trained decoder however, is not always beneficial to downstream performance in the case of SlotCon. More details can be found in Supplementary Material 5.
@@
Out-of-domain tasks: Table 4 depicts the fine-tuning performance of a model pre-trained on COCO to out-of-domain datasets. We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased. Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if only the encoder was transferred. We also observed that pre-trained models outperformed models trained with randomly initialized weights. This gap was more notable in REFUGE dataset (over 20%).
@@
Table 4: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights. We pre-trained SlotCon and its DeCon-SL adaptation with a ResNet-50 encoder and an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets. We fine-tuned on REFUGE and ISIC datasets under various dataset sizes. We averaged the fine-tuning mIoU results over three runs.
@@
4.2 Ablation studies
Table 5 presents an ablation study of the components in our DeCon-ML-L framework. The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, even when only the encoder is transferred, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision. This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
Table 5: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance. We averaged all fine-tuning results over three runs. $\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
Framework Improvement COCO Obj Det. COCO Inst. Seg.
AP $\Delta$ AP $\Delta$
SlotCon 40.81 - 36.80 -
+FPN Decoder (Multi Level Loss) 40.85 +0.04 36.83 +0.03
+Dropout = 0.5 41.10 +0.25 37.05 +0.22
+Enc-Dec Loss Weight ($\alpha$) Tuning (DeCon-ML-L) 41.18 +0.08 37.12 +0.07
Table 6 shows DeCon-ML performance with different values of encoder loss weight $\alpha$ and channel dropout. A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L. Hence, this value was chosen to ablate $\alpha$. An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own. For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
Table 6: Impact of the channel dropout probability and of the encoder loss weight $\alpha$ on DeCon. We pre-trained DeCon adaptations of SlotCon framework on COCO dataset using different parameters and evaluate the pre-trained encoder fine-tuning on downstream tasks that use the same decoder architecture. We averaged all fine-tuning results over three runs.
Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Frame work Deco der $\alpha$ Sem. Seg.
COCO COCO VOC City
AP AP mIoU mIoU
Dropout
DeCon-ML-L
FPN
0.50 0 40.85 36.83 - - - - -
0.25 40.95 36.97 - - - - -
0.50 41.10 37.05 - - - - -
0.75 41.02 37.05 - - - - -
$\alpha$
0.75 0.50 40.98 36.90
DeCon-SL (No Drop)
FCN
0.75 71.98 76.18
0.50 41.10 37.05 0.50 72.42 75.62
0.25 41.05 37.13 0.25 73.01 76.21
0 41.18 37.11 0 72.17 75.71